4. ULADisaster:4 A200-kgrocketindeepspacemoveswithavelocityof121ms^−1ˆi+38.0ms^−1ˆj. Suddenly, it explodes into three pieces, with the first (78kg) moving at −321ms^−1ˆi + 228 m s^−1 ˆj and the second (56 kg) moving at 16.0 m s^−1 ˆi − 88.0 m s^−1 ˆj. Find the velocity of the third piece.
Answer #1

let the mass of the third particle be m3

by the conservation of mass we have

m1 + m2 + m3 = 200

78 + 56 + m3 = 200

m3 = 66 kg

Let the x-component be Vx and y-component be Vy

Conserving the momentum in x-direction we have

200*121 = 78*-321 + 56*16 + 66vx

vx = 732.45 m/s

Conserving the momentum in y-direction we have

200*38 = 78*228 + 56* -88 + 66Vy

Vy = -79.63 m/s

so velocity of third piece is 732.45i - 79.63j